OpenStudy (anonymous):

if x=3 and y=-2 then y^4 - 4x?

OpenStudy (anonymous):

-2^4= 16 4*3=12 16-12=4

OpenStudy (anonymous):

-2^4 - 4(3) 16 - 12 4